OpenStudy (ryandane):

I am not sure about the answer but what is 69 and 69 isn’t it dinner for 4 ? O.O

OpenStudy (fortytherapper):

?

OpenStudy (unklerhaukus):

4d = 69+69 = 138 d = 138/4 = 34.5